Recent Maintenance – BFO

BFO operational status: up


There was a lightning strike near the station that caused some damage to the network and timing for seismic data to be lost.  The station operator switched network cables and restored normal operations at the station.


The North-South channels from the STS-1N seismometer were non-seismic since from day 2015,141-14:45 to 2015,142-18:10 due to a faulty connection to the seismometer. The channels were restored after the station operator reconnected the cable.


The station is transitioning from the Mk7 system to a Quanterra Q330HR system. Data is being acquired and will be sent to the DMC from all instruments. However, at present some of the instruments are still connected to the Mk7 system.


The vertical channels from the STS-1 seismometer are back in normal operating range after the station operator re-installed the STS-1 seismometer on day 2012,167-15:41.


The station operator reports problems with the STS-1V holding vacuum. The vertical channels have flat-lined intermittently since day 2012,155-20:45. The station operator is planning on re-installing the STS-1V seismometer.


The VSB channel from the Cambridge Invar wire strainmeter became non-functional after the Chilean quake on day 2010,059. This channel was restored after the station operator changed the silica gel for the instrument on day 2010,102.


Data is being withheld for the VGZ,WDI, UAN, and UAE channels. The UAN and UAE channels have been disconnected since about day 2009,098. They will be reconnected after the Askania pendula is relocated within the site.


Data quality for the VGZ channel improved as of about day 2007,231-0735, and for the WDO channel as of about day 2007,228-1840.


Data quality for the VGZ channel from the gravimeter and the WDO channel from the microbarograph is poor as of about day 204-06:28 for reasons yet to be determined. The station operator reported a power outage due to a lightning strike.


The E/W channels for the STS-1 seismometer are back after the station operator replaced the faulty RT303 board in the DAS.


The E/W channels for the STS-1 seismometer are out due to a faulty RT303 board in the DAS.


The response for the VAN and VAE channels from their original installation to the present has been corrected. The former response for these channels was incorrect.


Station was hit by lightning. Several buffer amplifiers got fried. Five hours later, the station operation is back to normal after the damages were repaired.


New EPROMs installed to correct DSP problem.


Station is operating normally.